Please verify that my son has solved the x in the two equations below and that he has correctly verified the solutions.
Equation 1
5(x-3) +(2(3-x) = 12
5x - 15 + 6 - 2x = 12
5x - 2x - 15 + 6 = 12
3x -9 = 12
3x -9 + 9 = 12 + 9
3x = 21
x = 7
Solution
5 (7-3) + 2 (3 -7 = 12
5(4) + 2 (-4) = 12
20 + -8 = 12
12 = 12
Equation 2:
18 - (x+3) - 2x = 4 + (x-5)
18 - x +3 - 2x = 4 + x-5
-x - 2x +18 + 3 = 4 + x-5
-3x +21 = 4 + x-5
-3x - x + 21 - 21 = x - x + 4 -5 -21
4x = -22
x = -5.5
Solution:
18 - 5.5 + 3 -2(-5.5) = 4 + -5.5 -5
18 - 2.5 + 11 = 4 + -5.5 - 5
18 - 11 - 2.5 = 4 + -5.5 - 5
4.5 = -6.5
Notice on the 2d solution, cannot get it to equal or balance - please verify if Equation 1 and solution is correct.
On 2d equation, please verify the correct equation and solution for it.
Thank you very much for your assistance.

Answer by greenestamps(13200)



If you are trying to check the solution to see that the equation balances, and it does not, then you know the solution you got is not right; that means there is something wrong with the work you did.

Here is an error I see...:

18 - (x+3) - 2x = 4 + (x-5)
18 - x +3 - 2x = 4 + x-5 No; this line is not equivalent to the previous line

There appear to be other errors later on in the work you show. So don't try to "fix" the work you did earlier. Instead, fix the error I pointed out and work the problem anew from there.

By the way.... Thanks for showing in detail the work you did. Despite instructions at the forum that people submitting questions should show the work they have done, only 1 in 100 (if that many) do.